OpenStudy (anonymous):

Use the Distributive Property and simplify the expression. 8 + 3(2x – 1) + 12

OpenStudy (anonymous):

I will give a medal to the best answer!! I really need this!

OpenStudy (teddyiswatshecallsme):

Do you know what the distributive property is? Like what it states.

OpenStudy (anonymous):

No :? It's a question my math teacher wants me to solve and I never got this before.

OpenStudy (teddyiswatshecallsme):

mmk.... Well, basically whenever you see a equation that has paranthesis, depending on the instructions and what they want to you do with the equation... In this case you need to simplify the expression 8 + 3 = 3(2x - 1) + 12 You would basically take the 3(2x - 1) and multiply 3 by everything that is in the parenthesis, so that'd be 3x 2x and 3 times 1, so that would be 6x - 3 so then the expression simplifies down to 8 + 3 = 6x - 3 + 12 now thats simplifying..... I could tell you, how to evaluate if thats what you need to do... what are the options? is it multiple choice?

OpenStudy (anonymous):

Heres the answers. A. 6x + 19 B. 5x + 17 C. 6x + 23 D. 6x + 17

OpenStudy (teddyiswatshecallsme):

Do you know how to take screenshots?